OpenStudy (whiskyryan):

Can someone show me why (1)/(2-sqrt3) = (2+sqrt3)/1 ?

OpenStudy (eliesaab):

Multiply up down by\[ 2+\sqrt 3 \]

OpenStudy (eliesaab):

\[ \frac 1 { 2-\sqrt 3}\frac {2+\sqrt 3}{2+\sqrt 3} \] What would you get?

OpenStudy (eliesaab):

The numerator would become \[ 2 + \sqrt3 \]

OpenStudy (eliesaab):

What would the denominator be?

OpenStudy (whiskyryan):

2^2-sqrt^2 = 4-3 = 1 ... Thanks!